# Resilient and Adaptive Replanning for Multi-Robot Target Tracking with Sensing and Communication Danger Zones

### v1

@@ IV-C3 Sensing and Communication Attacks
Consider a robot that is located in an overlapping region of sensing and communication danger zones and is thus subject to attacks from multiple sources. Without any measurement or communication ability, the mission of such a robot would be to perform single-agent planning to escape from the dangerous region as soon as possible. Its single-agent planning is done by solving an optimization similar to Eq. 12, except that its objective function is revised as:
@@
i.e., to escape from the dangerous region with minimum control effort. In addition, the weights for different danger zones are adaptive so that the robot can prioritize the zone(s) where it is already in.
@@ IV-D Estimation Update and Recover
As we have described in Sec. IV-C, if a robot is under communication attack, it will perform single-agent target tracking using its onboard sensing while attempting to escape from the danger region since it cannot pass messages to its teammates. In this case, the robot under attack and the rest of the team with normal communication run two separate EKFs to estimate the state of targets. As soon as the robot successfully escapes from the communication danger zone, it broadcasts its estimate to teammates. To merge estimates from the two Kalman filters, we use covariance intersection (CI). Let the estimate as well as covariance matrix from the robot under communication attack be $\hat{\mathbf{z}}_{1}$ and $\mathbf{\Sigma}_{1}$, the estimate and covariance matrix from the robots with normal communication be $\hat{\mathbf{z}}_{2}$ and $\mathbf{\Sigma}_{2}$. Then, CI performs the following operations to get the merged estimate $\hat{\mathbf{z}}_{m}$ and the merged covariance matrix $\mathbf{\Sigma}_{m}$:
@@
where $\omega$ is a parameter, and its value can be obtained by optimizing a particular norm [27]. From there, the robot which previously lost communication and the rest of its teammates becomes a single team once again. They can resume target state estimation using one standard EKF.
IV-E Adaptive Replanner
To further improve the performance, we propose a real-time adaptive replanning strategy that adjusts operations based on varying known conditions within danger zones. Here, each robot has its distinct weight parameters, and the adaptive replanner will tune the weight to adjust the strategy for each robot. For instance, the default weights for all the robots are initially prioritized for tracking performance. This will relax the constraints on robots’ trajectory, allowing them to take more risk and cut through the danger zones for better tracking performance. When an attack occurs on the robot, the adaptive planner will tune the weight to reduce the risk and the robot can stay further from the zones.
@@ V-A Performance Analysis
We conduct several simulation experiments involving sensing attacks, communication attacks, and a combination of both to investigate the performance of our framework. For brevity, the sensing danger zone is referred to as the Type I zone, while the communication danger zone is denoted as the Type II zone. More scenarios are available in our open-source code: https://github.com/Zhourobotics/resilient-target-tracking.git.
V-A1 Tracking under Senser Attack
We demonstrate the scenario of two robots tracking two targets with an unknown sensing danger zone. We set $\delta_{1}$ as 0.1 and the attack frequency of the sensing danger zone as 1 Hz. We perform sampling every 0.1 seconds with a maximum of 300 steps. The trajectories and attack status under two different $\epsilon$ values, which impose different levels of risk requirements, are shown in Fig. 2. Initially, robots have no information about the sensing danger zone and focus on tracking the targets. Then Robot 0 encounters an attack from the danger zone and immediately broadcasts information about this danger zone to Robot 1. Notably, the robots can still be attacked even after they have discovered a danger zone, which depends on the risk requirements. For instance, compared with (b.1) - (b.3) in Fig. 2, robots suffer from more attacks in (a.1) - (a.3) due to a larger risk bound $\epsilon$, but stay closer to target for better tracking quality.
Figure 2: Tracking with a sensing danger zone. The mean position for the center of the danger zone is at (0.1, 0) with a covariance of 0.3. Figure (a.1) - (a.3) are the trajectories when $\epsilon=0.2$, and for (b.1) - (b.3), $\epsilon=0.1$.
Figure 3: Tracking with two communication danger zones. The first danger zone center point has its mean [-0.3, 0.0] and covariance $\Sigma=0.3$, the second one centered at [1.0, 0.2] with $\Sigma=0.2$. We set $\epsilon=0.9$ for (a.1) - (a.3), and $\epsilon=0.95$ for (b.1) - (b.3).
Figure 4: Snapshot of two robots tracking two targets under multiple attacks. The figure from left to right demonstrates the whole process from initially unknown to known of the danger zones. Robot 0 experiences a sensing attack in Type I danger zone 1, and then successfully recovers. Subsequently, Robot 1 is subjected to communication attacks in a communication danger zone. This process simulates multiple attacks and recoveries across both types of danger zones while gathering critical information about each zone.
@@ V-A2 Tracking under Communication Attack
We use a similar setup for communication attacks to illustrate how the model differs from that used for sensing danger zones and to highlight the corresponding variations in robot behaviors. As shown in Fig. 3, we consider two robots tracking four targets, and each robot is assigned to track two targets accordingly. When robots are close to the communication danger zone, Robot 1 gets attacked and loses its communication link to Robot 0. Robot 1 prioritizes escaping from the danger zone and recovering from the communication failure. Then, it shares information about the danger zone with Robot 0 such that the latter avoids the danger zone precautiously.
@@ V-A3 Tracking under Multiple Attacks
We consider a more complicated scenario involving two robots tracking two targets through three sensing danger zones and two communication danger zones in the environment. The sampling time is 0.2 s, with a maximum of 300 steps. As shown in Fig. 4, the two robots can efficiently track the targets while balancing the importance of tracking quality and safety. When the robot passes through the overlapping area of two different types of danger zones, a more aggressive strategy enables it to navigate the areas and successfully complete its tracking task.
V-B Ablation Study for Adaptiveness
We perform an ablation study comparing two methods: one without adaptiveness (denoted as "w/o adaptiveness") and one with adaptiveness ("w/ adaptiveness") in three types of scenarios described above. We evaluate the tracking performance by its average trace, the total number of attacks on all robots, and trajectory length. The result is shown in Tab. I. In all scenarios, the adaptive method leads to lower traces, indicating that the robots can efficiently track targets in the environment when they employ adaptiveness in their planning. In addition, the adaptive strategy leads the robots into more dangerous areas, potentially because they prioritize tracking quality over safety and energy in these cases.
TABLE I: Evaluation of Adaptiveness Performance
Scenarios Methods
Accumulated
Trace
Total
Attacks
Trajectory
Length (m)
Type I Zones w/o adaptiveness 16.47 61 12.66
w/ adaptiveness 15.70 75 13.57
Type II Zones w/o adaptiveness 37.80 61 9.55
w/ adaptiveness 30.89 85 9.54
Muti-type Zones w/o adaptiveness 28.74 122 17.34
w/ adaptiveness 26.45 149 18.41
V-C Comparison with Single-Agent Planner
@@
In the single-agent planner framework, when a robot encounters attacks from danger zones, it cannot share information with the rest of the team, leading to delayed avoidance of danger zones. Our proposed resilient planner framework allows a robot to share information after recovery to improve collaborative tracking performance.
We benchmark our method against sequential single-agent resilient tracking to validate the effectiveness of our proposed framework. The trajectory of individual tracking and our proposed tracking are shown in Fig. 5.
The robots with single-agent planners clustered in the center due to the lack of communication and, therefore, have a worse tracking performance.
@@
Figure 5: Simulation of two robots tracking two targets under multiple attacks with (a) proposed centralized framework and (b) single-agent planners.
V-D Hardware Experiments
We validate the robustness of our proposed framework through several hardware experiments in various initially unknown hazardous environments. We apply a similar experiment setting as in our prior work [11] with Crazyflie drones [28] as trackers and Yahboom ROSMASTER X3 ground robots [29] as targets. The scenario of a single danger zone is shown in Fig. 6, and that with multiple zones is shown in Fig. 1. The robots initially track the targets closely but switch to resilient behavior when one of them is attacked and identifies the presence of a danger zone. In the case of a sensing danger zone, all robots simultaneously escape the area as the robots can share the information immediately. However, in a communication danger zone, the response is sequential, with one robot escaping and sharing the information, followed by the second robot avoiding the zone after the first one escapes.
@@ VI Conclusion
In this paper, we propose a resilient and adaptive multi-robot target tracking architecture under unknown danger zones. We model the danger zones as risk fields with attack and recovery models to simulate different failure scenarios. We solve the problem with different optimization formulations and propose an adaptive replanner for robots to adjust the priority of task goals. In the future, we will extend the resilient framework to more complex and realistic scenarios, including real-time perception, limited resources, dynamic danger zones, and adversarial targets.

### v2

@@ IV-C3 Sensing and Communication Attacks
Consider a robot that is located in an overlapping region of sensing and communication danger zones and is thus subject to attacks from multiple sources. Without any measurement or communication ability, the mission of such a robot would be to perform single-robot planning to escape from the dangerous region as soon as possible. Its single-robot planning is done by solving an optimization similar to Eq. 12, except that its objective function is revised as:
@@
i.e., to escape from the dangerous region with minimum control effort. In addition, the weights for different danger zones can be adjusted so that the robot can prioritize the zone(s) which it is already in.
@@ IV-D Estimation Update and Recover
As we have described in Sec. IV-C, if a robot is under communication attack, it will perform single-robot target tracking using its onboard sensing while attempting to escape from the danger region since it is not part of the communication network and thus cannot exchange measurements with its teammates. In this case, the robot under attack and the rest of the team within the communication network run two separate EKFs to estimate the state of the targets. As soon as the robot successfully escapes from the communication danger zone and regains communication, it broadcasts its estimate to teammates in the communication network. To merge estimates from the two EKFs, we use covariance intersection (CI). Let the estimation as well as covariance matrix from the robot under communication attack be $\hat{\mathbf{z}}_{1}$ and $\mathbf{\Sigma}_{1}$, the estimation and covariance matrix from the robots within the communication networks be $\hat{\mathbf{z}}_{2}$ and $\mathbf{\Sigma}_{2}$. Then, CI performs the following operations to get the merged estimate $\hat{\mathbf{z}}_{m}$ and the merged covariance matrix $\mathbf{\Sigma}_{m}$:
@@
where $\omega$ is a parameter, and its value can be obtained by optimizing a particular norm [30]. After CI, all robots have the same estimate of targets, i.e., ($\hat{\mathbf{z}}_{m}$, $\mathbf{\Sigma}_{m}$).
@@ V-A Performance Analysis
We conduct a series of simulation experiments involving sensing attacks, communication attacks, and their combination to evaluate the performance of our resilient target tracking framework. Each robot can observe and track multiple targets. Our approach is compared to a vanilla target tracking framework that lacks resilient recovery mechanisms against attacks. More scenarios are available in our open-source code: https://github.com/Zhourobotics/resilient-target-tracking.git.
V-A1 Tracking under Sensing Attack
We demonstrate the scenario of two robots, $\mathit{R}_{0}$ and $\mathit{R}_{1}$, tracking two targets, $\mathit{T}_{0}$ and $\mathit{T}_{1}$, with an unknown sensing danger zone, $\mathit{C}_{0}$. We set $\delta_{1}$ as 0.1 and the attack frequency of the sensing danger zone as 1 Hz. We perform sampling every 0.1 seconds with a maximum of 300 steps. The trajectories and attack status from our resilient tracking and vanilla tracking without resiliency to attacks are shown in Fig. 2. Initially, robots have no information about the sensing danger zone and focus on tracking the targets. Then $\mathit{R}_{1}$ encounters an attack from the danger zone and immediately broadcasts information about this danger zone to $\mathit{R}_{0}$. Notably, the robots can still be attacked even after they have discovered a danger zone. The vanilla tracking method suffers from sensing attacks as shown in (b.1) - (b.3) from Fig. 2; the two robots become immobilized when both of them lose their measurements and cannot continue the target tracking mission. On the other hand, our resilient framework in (a.1) - (a.3) can efficiently identify, broadcast, and recover from the sensing attack. Fig. 3 demonstrates the mean tracking error and trace of the covariance matrix over ten tests for both the resilient and vanilla tracking methods. Our proposed method maintains high tracking accuracy even in the presence of sensor attacks, whereas the vanilla tracking method fails, resulting in error and trace explosion.
Figure 2: Tracking with a sensing danger zone. The mean position for the center of the danger zone is at [0.1, 0] with a covariance of 0.3. Figures (a.1) - (a.3) are the trajectories with resilient tracking, and (b.1) - (b.3) are those with the vanilla tracking.
Figure 3: Performance of the proposed resilient tracking (blue) versus vanilla tracking (red) under the conditions in Fig. 2, shown in MSE (a) and the trace of the covariance matrix (b) of the targets’ state estimations. The solid line denotes mean values with shaded regions as the $\pm$1 standard deviation.
@@ V-A2 Tracking under Communication Attack
We use a similar setup for the communication danger zone to illustrate the performance gap between our resilient tracking and the vanilla tracking with communication attacks. As shown in Fig. 4, we consider two robots tracking four targets with different parameter setups, and each robot is assigned to track two targets accordingly. When robots are close to the communication danger zone, $\mathit{R}_{0}$ gets attacked and loses its communication link to $\mathit{R}_{1}$. $\mathit{R}_{0}$ prioritizes escaping from the danger zone and recovering from the communication attack. Then, it shares information about the danger zone with $\mathit{R}_{1}$ so that the latter avoids the danger zone. However, for the vanilla tracking, the robots do not share information, and $\mathit{R}_{1}$ still tries to enter the danger zone after $\mathit{R}_{0}$ recovers from the communication attack. Fig. 5 shows similar results that the vanilla tracking performs poorly when receiving communication attacks from the environment compared with our resilient tracking.
Figure 4: Tracking with two communication danger zones. The first danger zone center point has its mean [-0.3, 0.0] and covariance $\Sigma=0.3$, the second one centered at [1.0, 0.2] with $\Sigma=0.2$. (a.1) - (a.3) show the trajectories from the resilient tracking, and (b.1) - (b.3) show the trajectories from the vanilla tracking.
Figure 5: Performance of the proposed resilient tracking (blue) versus vanilla tracking (red) under the conditions in Fig. 4, shown in MSE (a) and the trace of the covariance matrix (b) of the targets’ state estimations.
@@ V-A3 Tracking under Multiple Attacks
(a) Resilient tracking
(b) Vanilla tracking
Figure 6: Snapshot of two robots tracking two targets with three sensing and two communication danger zones. The sampling time is 0.2 s, with a maximum of 300 steps. Figures from left to right demonstrate the process of discovering the unknown danger zones over time. The robot $\mathit{R}_{0}$ experiences a sensing attack in sensing danger zone 1, and then successfully recovers. Subsequently, $\mathit{R}_{1}$ is subjected to communication attacks in a communication danger zone. This process simulates multiple attacks and recoveries across both types of danger zones while gathering critical information about each zone.
We consider a more complicated scenario involving two robots tracking two targets with multiple sensing and communication danger zones in the environment. As shown in Fig. 6, the two robots can efficiently track the targets while balancing the importance of tracking quality and safety. The danger zones are gradually revealed to the team as the individual robots receive attacks and recover. When the robot passes through the overlapping area of two different types of danger zones, a more aggressive strategy enables it to navigate the areas and successfully complete its tracking task. For vanilla tracking, when both robots lose communication abilities, they converge toward the center to maximize individual target coverage. In contrast, the resilient method enables robots to avoid the danger zone after revealing it. However, this central convergence in the vanilla tracking increases the risk of communication attacks. Fig. 7 shows that the tracking quality from the vanilla tracking suffers from attacks. Although both robots escaped from the attacks towards the middle of the tracking process, and the tracking quality improved, the following attacks deteriorated the tracking quality later in the process. On the other hand, our proposed resilient tracking framework provides consistent tracking performance under a sequence of attacks.
Figure 7: Performance of the proposed resilient tracking (blue) versus vanilla tracking (red) under the conditions in Fig. 6, shown in MSE (a) and the trace of the covariance matrix (b) of the targets’ state estimations.
V-B Comparison with individual resilient tracking
@@
With individual resilient tracking, when a robot encounters an attack in a danger zone, it escapes but does not share information with the rest of the team, causing delayed avoidance of danger zones for other robots. In contrast, our proposed resilient tracking enables a robot to share information after recovery, enhancing collaborative tracking performance.
We benchmark our collaborative resilient tracking against individual resilient tracking to validate the effectiveness of our proposed framework, as shown in Fig. 8. In the case of individual resilient tracking, robots cluster in the center due to the lack of information exchange about revealed danger zones, leading to poor tracking performance. In contrast, with our collaborative resilient tracking, robots share knowledge of danger zones after recovery, allowing them to maintain strong tracking performance throughout the process.
@@
Figure 8: Simulation of two robots tracking two targets under sensing and communication attacks with (a) the proposed collaborative resilient tracking and (b) the individual resilient tracking.
@@
V-C Hardware Experiments
We validate the resiliency of our proposed framework through multiple hardware experiments conducted in various initially unknown hazardous environments. Following a similar experimental setup as in our prior work [13], we use Crazyflie drones [31] as trackers and Yahboom ROSMASTER X3 ground robots [32] as targets. The scenario involving a single danger zone is illustrated in Fig. 9, while the case with multiple danger zones is shown in Fig. 1. In both scenarios, our resilient tracking framework enables robots to resiliently react and recover when one of them is attacked and identifies a danger zone. In the case of a sensing danger zone, all robots receive the danger zone information simultaneously and immediately escape, as they can share information in real time. However, in a communication danger zone, the response is sequential, where one robot escapes and recovers first before sharing the information, allowing the next robot to avoid the zone accordingly. This strategy ensures effective tracking performance even in the presence of attacks and environmental uncertainties.
@@ VI Conclusion
We present a resilient multi-robot multi-target tracking architecture under unknown sensing and communication danger zones. We address this multi-robot planning problem through an optimization formulation that allows robots to adjust the prioritization of task goals, considering factors such as tracking performance, safety, and efficiency. Simulations and real robot experiments show the resiliency of our proposed framework when encountering sensing and/or communication attacks. In the future, we will extend the resilient framework to more complex and realistic scenarios, including real-time perception, restricted resources, limited sensing and communication ranges, dynamic danger zones, and adversarial targets.